OpenStudy (tywower):

Line L has equation 2x - 3y = 5. Line M passes through the point (2, -10) and is perpendicular to line L. Determine the equation for line M.

OpenStudy (tywower):

A) 2x + 3y = -14 Eliminate B) 2x - 3y = -26 C) 3x + 2y = -14 D) 3x - 2y = -26

OpenStudy (tywower):

@TheSmartOne ?

OpenStudy (tywower):

@Nnesha

OpenStudy (tywower):

@misty1212

TheSmartOne (thesmartone):

What is the relation between the slopes of perpendicular lines?

OpenStudy (gorv):

what is the relation between 2 perpendicular lines ??

OpenStudy (tywower):

idk

OpenStudy (gorv):

product of slope of two perpendicular line = -1

OpenStudy (tywower):

@cleetus779

OpenStudy (tywower):

@triciaal

OpenStudy (gorv):

2x-3y=5 can u tell the slope of this line ??

OpenStudy (tywower):

2

TheSmartOne (thesmartone):

Make that into the form of y=mx+b and m will be the slope

OpenStudy (gorv):

first convert it in point slope form

OpenStudy (tywower):

I don't feel like it

TheSmartOne (thesmartone):

@tywower If you are only interested in the final answer, then I'm sorry because I'm here to guide people to the answer not give them the answer.

TheSmartOne (thesmartone):

And also, it is against the Code of Conduct to give answers. http://openstudy.com/code-of-conduct

OpenStudy (tywower):

I KNOW I KNOW IT's just that I have no memory ON HOW TO DO IT or what point slope is

TheSmartOne (thesmartone):

I told you what point slope form is. It's in the form of \(\sf y=mx+b\) So, this is your question: 2x-3y=5 Just take 2x to the other side, and then divide by -3 to both sides.

OpenStudy (tywower):

@Preetha

TheSmartOne (thesmartone):

Just tagging other users will not help get you an answer.

TheSmartOne (thesmartone):

I'm still waiting right here for you to cooperate. If you don't want to, I'll be on my way to check my 100+ notifications.

OpenStudy (triciaal):

do you still need help?